# Climate Physical Risk Assessment in Asset Management

### v1

@@ 5 Evaluation of physical risk impacts
In our framework, the dynamics of the asset values remain consistent with Vasicek’s risk-neutral paradigm, since the underlying stochastic evolution is not altered. The crucial difference, however, lies in debt management: instead of a single maturity, a continuous (rolling) refinancing mechanism is adopted, which continually extends the maturity. This approach better aligns with standard corporate practice, as firms typically refinance their debt over time rather than adhering to a single fixed maturity. Consequently, this rolling structure leads to an equity payoff that differs from the classical formulation, where maturity is predetermined. As a result, while preserving internal consistency with the risk-neutral framework, the measure of the expected loss may differ from the one determined under the traditional Vasicek model.
[…]
At a 1-year horizon, all indexes show similar increases in $\Delta L$ and $\Delta VaR$, with VaR growing across different confidence levels. Over longer horizons, the differences become more pronounced, and both $\Delta L$ and $\Delta VaR$ increase over time for all indexes. Notably, at longer time horizons, such as 20 years, $\Delta VaR$ does not increase with higher quantiles. This is because, at extended time horizons, the equity value, which is modeled as a European Call option on the asset value, is very close to zero for extreme quantiles of both processes with and without jumps. Therefore, there is no significant difference for extreme quantiles, while we still see an impact on the $\Delta L$. We point out that, for an asset manager perspective, the relevant time-horizons is from one to five years for which the $\Delta VaR$ increases over quantiles. In this analysis, we assume no correlation between the Lévy drivers $L_{k}(t)$ that model the climate shocks for each cluster. For robustness check, in Appendix Appendix A: Correlation Robustness, we focus on the extreme correlation case, i.e. when the correlation between the Lévy drivers is the highest possible. We do not find significant differences in the case of correlated jump processes.
@@
Table 6: $\Delta L$, $\mathbf{\Delta VaR_{90\%}}$, $\mathbf{\Delta VaR_{95\%}}$, and $\mathbf{\Delta VaR_{99\%}}$ for MSCI World (A), MSCI World ESG Leaders, and MSCI World Climate Paris Aligned indexes over various time horizons. Results are derived from simulated dynamics under the risk-neutral measure.
@@ C Low Medium High Extreme
Medium 0.53% 4.45% 2.04% 0.62%
Table 7: Clustering results for MSCI World (A), MSCI World ESG Leaders (B), and MSCI World Paris Aligned (C) respectively. Columns represent asset intensity clusters, while rows indicate the regional vulnerability cluster of belonging.
@@
[…]
@@ 6 Conclusions & Policy implications
In this paper, we introduce a novel model for quantifying the impact of physical climate risk on an equity portfolio, utilizing a firm value model (FVM) approach.
We introduce a new and easily implementable modeling framework for estimating physical climate risk at the firm level, which can be seamlessly integrated into financial institutions’ decision-making processes and customized for various climate risk metrics. The approach extends the multivariate Vasicek model in (2) by incorporating negative jumps to account for climate-related losses. An explicit formula for the stressed equity value, based on the characteristic function of the jump process, is provided in (7).
We calibrate the model using estimated climate risk losses in conjunction with widely available balance sheet data. Companies are clustered according to their physical risk profiles by leveraging their geographical vulnerability and asset intensity scores. This results in two clusters based on asset vulnerability and four clusters corresponding to increasing levels of asset intensity (cf. Figures 24 and Tables 12). The FVM parameters are then calibrated for each cluster by applying the two-step procedure outlined in Section (4.3).
We apply the calibrated model to compute the differences in Value-at-Risk (VaR) and expected loss between a baseline scenario and a climate-stressed scenario across various portfolios (see Table 3 and Figure 5). Building on the VaR computation, we propose an add-on to the portfolio VaR that explicitly accounts for physical climate risk (see Figure 6). This framework highlights the heterogeneous exposure to climate risk across sectors such as utilities, which exhibit higher vulnerability.
From a policy perspective, our framework facilitates the calculation of a VaR add-on for any equity portfolio, given knowledge of each stock’s geographical location, sector and widely available balance sheet data. This enables regulators to define standardized VaR add-ons that could be potentially integrated into the Basel III standard approach, without requiring supervised entities to implement or replicate the model independently.
@@ Acknowledgments
We thank N. Bartolini, R. Baviera, R. Ceretti, U. Cherubini, P. Manzoni, E. Sala, T. Vargiolu for the fruitful conversations on the topic and all the participants to the III International Fintech Conference 2025, the Energy Finance X Conference and the CLIFIRIUM 2025 Conference. Michele Azzone and Davide Stocco are members of the Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INdAM).
Appendix A: Correlation Robustness
@@
Appendix B: Climate Stress Scenario Simulation
In this Appendix, we explain in detail the algorithm for simulating the climate stressed scenarios and computing losses on various portfolios. Moreover, we report additional robustness checks regarding the computation of losses when the investor does not neglect climate risk.
@@
1:Initialize $Loss$ and $Loss_{shock}$

### v2

@@ 5 Evaluation of physical risk impacts
In our framework, the dynamics of the asset values remain consistent with Vasicek’s risk-neutral paradigm, since the underlying stochastic evolution is not altered. The crucial difference, however, lies in debt management: instead of a single maturity, a continuous (rolling) refinancing mechanism is adopted, which continually extends the maturity. This approach better aligns with standard corporate practice. Corporates commonly manage debt rollover through a process known as debt refinancing, which involves issuing new debt to repay maturing obligations. This practice allows firms to maintain liquidity, optimize capital structure, and potentially benefit from favorable interest rates (Valenzuela 2016). Consequently, this rolling structure leads to an equity payoff that differs from the classical formulation, where maturity is predetermined. As a result, while preserving internal consistency with the risk-neutral framework, the measure of the expected loss may differ from the one determined under the traditional Vasicek model.  [^11] We have also repeated the analysis under the assumption of fixed debt maturity. The results remain consistent with the baseline case, although we observe higher expected losses. This increase is attributable to the loss of the time-value component of the equity, which is modeled as a European call option on the firm’s asset value.
We analyze the simulation results for the $\Delta L=\mathbb{E}[\tilde{L}(t)]-\mathbb{E}[{L(t)}]$ and the ${\Delta VaR}=VaR\;\tilde{L}(t)-VaR\;L(t)\;$ at 90%, 95%, 99% level respectively for different time-horizons t=1, 5, 10, 20t=1,\,5,\,10,\,20 years and for different indexes. The data concerning the composition of the indexes is based on information as of 31 December 2023. In Table 6, we present the results for three relevant indexes: the MSCI World index which represents a broad benchmark covering developed markets, serving as a standard reference for global equity performances; the MSCI World ESG Leaders and Climate Paris Aligned indexes which offer specialized versions, with the ESG Leaders index focusing on companies excelling in environmental, social, and governance practices, and the Climate Paris Aligned index prioritizing firms aligned with the Paris Agreement’s climate targets. The results are very similar because the three indexes exhibit comparable asset intensity and regional vulnerability profiles, as shown in Table 7.
At a 1-year horizon, all indexes show similar increases in $\Delta L$ and $\Delta VaR$, with VaR growing across different confidence levels. Over longer horizons, the differences become more pronounced, and both $\Delta L$ and $\Delta VaR$ increase over time for all indexes. Notably, at longer time horizons, such as 20 years, $\Delta VaR$ does not increase with higher quantiles. This is because, at extended time horizons, the equity value, which is modeled as a European Call option on the asset value, is very close to zero for extreme quantiles of both processes with and without jumps. Therefore, there is no significant difference for extreme quantiles, while we still see an impact on the $\Delta L$. We point out that, for an asset manager perspective, the relevant time-horizons is from one to five years for which the $\Delta VaR$ increases over quantiles. In this analysis, we assume no correlation between the Lévy drivers $L_{k}(t)$ that model the climate shocks for each cluster. For robustness check, in Appendix B, we focus on the extreme correlation case, i.e. when the correlation between the Lévy drivers is the highest possible. We do not find significant differences in the case of correlated jump processes.
@@
Table 6: $\Delta L$, $\Delta VaR_{90\%}$}, $\Delta VaR_{95\%}$}, and $\Delta VaR_{99\%}$} for MSCI World (A), MSCI World ESG Leaders, and MSCI World Climate Paris Aligned indexes over various time horizons. Results are derived from simulated dynamics under the risk-neutral measure.
@@ C Low Medium High Extreme
Medium-High 0.53% 4.45% 2.04% 0.62%
Table 7: Clustering results for MSCI World (A), MSCI World ESG Leaders (B), and MSCI World Paris Aligned (C) respectively. The rows correspond to vulnerability clusters, while the columns represent intensity clusters.
@@
$\%\ \text{Add-on (t)}=\left(\frac{{VaR\;\tilde{L}(t)}}{VaR\;L(t)}-1\right)\cdot 100\;\;.$
@@
In Appendix C, we explain in detail the algorithm for simulating the climate-stressed scenarios and computing losses on various portfolios. Moreover, we report additional robustness checks regarding the computation of losses when the investor does not neglect climate risk.
@@ 6 Conclusions & Policy implications
In this paper, we introduce a novel and easily implementable model framework for quantifying the impact of physical climate risk on an equity portfolio, utilizing an FVM approach. The framework is designed to be seamlessly integrated into the decision-making processes of financial institutions and can be tailored to support a variety of climate risk metrics, depending on the specific regulatory or internal risk management objectives.
Our approach extends the multivariate Vasicek model in (2) by incorporating negative jumps to capture the sudden and severe financial impacts associated with extreme climate events. An explicit formula for the stressed equity value, based on the characteristic function of the jump process, is provided in (7). These negative jumps are introduced to model abrupt losses in firm value that may result from physical damages or operational disruptions. For instance, a wildfire might destroy a firm’s warehouse, leading to immediate asset write-downs, while an electricity shortage caused by a heatwave could halt production and severely impair revenue streams. By explicitly accounting for such climate-induced jumps, the proposed framework enhances the realism and robustness of a traditional market risk model, making it more suitable for analyzing portfolio physical risk. We underline that the model is constructed such that climate-related losses directly impact the firm’s asset value, thereby ensuring that the transmission channel to equity prices is explicitly and transparently reproduced. This direct linkage allows for a more accurate assessment of how extreme physical climate events propagate through to market valuations.
We calibrate the model using estimated climate risk losses in conjunction with widely available balance sheet data, such as total assets and asset intensity. To account for firm-level heterogeneity in exposure to physical climate risk, we classify companies into clusters based on two key dimensions: geographical vulnerability and asset intensity. Specifically, we construct two vulnerability clusters—distinguishing firms operating in high- versus low-risk regions—and four clusters representing increasing levels of asset intensity (cf. Figures 24 and Tables 12). We highlight that the cluster associated with high vulnerability scores corresponds almost perfectly to countries in the Global South. This empirical finding reflects the well-documented asymmetry in the geographic distribution of climate-related physical risks, with developing economies being disproportionately affected by extreme weather events, rising temperatures, and sea-level rise.
The model parameters are then estimated for each cluster using the two-step procedure described in Section 4.3, allowing for cluster-specific dynamics in response to climate shocks. The decision to group firms by both asset intensity and geographical exposure is motivated by the expectation that companies located in similar regions and with similar shares of physical assets will be affected in comparable ways by extreme climate events. This clustering strategy enhances the model’s ability to capture intra- and inter-sectoral dependencies, which is essential for producing more realistic portfolio-level risk assessments.
We apply the calibrated model to compute the differences in Value-at-Risk (VaR) and expected loss between a baseline scenario and a climate-stressed scenario across various portfolios (see Table 3 and Figure 5). We analyze simulation results for expected losses ($\Delta L$) and changes in Value-at-Risk ($\Delta VaR$) at the 90%, 95%, and 99% confidence levels over time horizons of 1, 5, 10, and 20 years, across three major indexes: MSCI World, MSCI World ESG Leaders, and MSCI Climate Paris Aligned. These indexes exhibit similar asset intensity and regional vulnerability profiles, resulting in comparable climate risk impacts. At short horizons (e.g., 1 year), $\Delta L$ and $\Delta VaR$ show consistent increases across all indexes and quantiles. No significant differences were observed between the MSCI World index and its ESG-aligned versions, highlighting that investing in ESG assets does not necessarily offer protection against the financial impacts of physical climate risk. [^13] For a review of ESG scores see e.g. Ferro et al. (2025). Building on the VaR computation, we propose an add-on to the portfolio VaR that explicitly accounts for physical climate risk (see Figure 6). This framework highlights the heterogeneous exposure to climate risk across sectors such as utilities, which exhibit higher vulnerability, with the highest VaR add-on of 86%. Our finding is motivated analyzing the firms that are included in the MSCI utility index (Figure 7). We observe that the majority of firms within the index fall into the extreme intensity category. These firms typically maintain large, immobile physical assets—such as power plants, transmission networks, and water systems—which are particularly vulnerable to climate-induced disruptions.
From a policy perspective, our framework enables the computation of a Value-at-Risk (VaR) add-on for any equity portfolio, provided that the geographical location, sector classification, and widely available balance sheet data of the constituent stocks are known. This facilitates the definition of standardized VaR add-ons by regulatory authorities, which could potentially be incorporated into the Basel standardized approach. Importantly, this would not require supervised entities to independently implement or replicate the underlying climate risk model. As an illustrative example, regulators could compute VaR add-ons using representative portfolios composed of the most asset-intensive stocks available in a given market (e.g., the U.S. equity market). The resulting add-ons could serve as conservative benchmarks, applicable to all portfolios that are exclusively exposed to the same geographical region. Moreover, we highlight that our calibration approach facilitates regulatory and policy applications, as it enables authorities to derive generalized insights on climate vulnerability patterns and apply them to supervisory stress testing exercises. It also reduces model complexity for institutions by allowing them to apply pre-calibrated cluster-level parameters based on simple, observable firm characteristics.
@@
Limitations and Further developments
While the proposed framework provides a tractable and flexible approach to modelling physical climate market risk, it also presents several limitations that open avenues for future research. A primary constraint lies in the geographical treatment of asset exposure: we consider only the location of firms’ headquarters, neglecting the precise geographical distribution of individual assets. This simplification may lead to misclassification of climate exposure, especially for multinational firms with physical assets spread across diverse regions. A natural extension, subject to data availability, would be to assign assets to different climate clusters based on their specific geographical positions or to determine a firm’s location based on where the majority of its physical assets are located.
Another modelling limitation stems from the assumption that climate shocks follow a Lévy process. While Lévy processes offer analytical tractability and encompass a broad class of jump processes, they are characterized by independent and stationary increments. Such assumptions may be too simplistic for modelling climate phenomena, which tend to be non-stationary and exhibit temporal dependencies. To better capture these features, the framework could be extended by incorporating additive processes - processes with independent but non-stationary increments (see, e.g., Azzone and Baviera (2022)) - or Hawkes processes - self-exciting processes in which past events increase the likelihood of future shocks (see, e.g., Hawkes (2018)). These alternatives could provide a more realistic representation of the complex, path-dependent nature of climate dynamics.
@@ Acknowledgments
We thank N. Bartolini, R. Baviera, G. Bressan, R. Ceretti, U. Cherubini, F. Maglione, P. Manzoni, E. Sala, A. Sbuelz, L. Taschini, T. Vargiolu for the fruitful conversations on the topic and all the participants to the III International Fintech Conference 2025, the Energy Finance X Conference and the CLIFIRIUM 2025 Conference. Michele Azzone and Davide Stocco are members of the Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INdAM). The research of Davide Stocco was carried out in the framework of “Energy for Climate” interdisciplinary research center and was supported financially by the “Decarbonize energy” program of the Institut Polytechnique de Paris.
Appendix A: Dataset descriptive statistics
In Table 8, we report a snapshot of 16 firms in the dataset. For each of the 16 randomly extracted firms we report the ISIN code, the Country, the NACE sector code, the MSCI index they are part of, the Vulnerability and the Intensity cluster the total debt (in billion local currency) the TTM (in years) and the market capitalization (in billion local currency).
ISIN Country NACE Index Vulnerability Intensity Debt TTM Mkt Cap
CA82509L1076 Canada G MSCI USA Low Low 0.9 2.5 99.1
FR0000039299 France M MSCI Europe Low Low 7.8 4.5 16.5
PHY806761029 Philippines G MSCI Emerging Asia Medium-High Low 541.1 3.8 1050.4
SA15GG53GHH3 Saudi Arabia M MSCI EMEA Medium Low 0.1 6.4 63.3
US7443201022 USA K MSCI USA Low Medium 21.1 8.9 38.2
US1729674242 USA K MSCI USA Low Medium 501.8 5.1 99.6
BRGGBRACNPR8 Brazil C MSCI Emerging America Medium-High Medium 13.6 9.1 27.3
JP3818000006 Japan J MSCI Pacific Medium-High Medium 211.2 0.5 4142.4
US9078181081 USA H MSCI USA Low High 33.3 14.6 151
GRS419003009 Greece R MSCI EMEA Low High 0.8 6.4 5.4
KYG4712E1035 China Q MSCI Emerging Asia Medium-High High 1.3 3.8 21.8
BRRDORACNOR8 Brazil Q MSCI Emerging America Medium-High High 37.2 8.2 64.8
AN8068571086 USA B MSCI USA Low Extreme 12.2 8 73.8
US5529531015 USA I.55 MSCI USA Low Extreme 8.9 3.7 17.2
SG1U68934629 Singapore D MSCI Pacific Medium-High Extreme 10.4 5.5 12.4
CNE [^00] T18 China B MSCI Emerging Asia Medium-High Extreme 7.8 3.8 69.9
Table 8: Snapshot of firms in the dataset. For the 16 randomly extracted firms we report the ISIN code, the Country, the NACE sector code, the MSCI index they are part of, the Vulnerability and the Intensity cluster the total debt (in billion local currency) the TTM (in years) and the market capitalization (in billion local currency).
In Table 9 we report descriptive statistics (mean, standard deviation, median, 5th and 95th percentiles) of Asset Intensity and Geographical Vulnerability computed at the country level. The reported moments are therefore obtained by averaging across countries. In addition, we provide descriptive statistics for the number of firms per country and for the number of firms per sector, to highlight the distribution of firms in our dataset across geographical and sectoral dimensions.
Mean Standard Deviation Median Q0.05 Q0.95
Asset Intensity per country 1.42 3.29 0.82 0.09 3.89
Vulnerability per country 0.44 0.23 0.45 0.11 0.92
# Firms per country 148.99 344.66 14.50 1 521.40
# Firms per sector 16.26 36.21 6 1 62.05
Table 9: We report descriptive statistics (mean, standard deviation, median, 5th and 95th percentiles) of Asset Intensity and Geographical Vulnerability computed at the country level. The reported moments are therefore obtained by averaging across countries. In addition, we provide descriptive statistics for the number of firms per country and for the number of firms per sector, to highlight the distribution of firms in our dataset across geographical and sectoral dimensions.
Appendix B: Correlation Robustness
@@
Appendix C: Climate Stress Scenario Simulation
In this Appendix, we explain in detail the algorithm for simulating the climate-stressed scenarios and computing losses on various portfolios. Moreover, we report additional robustness checks regarding the computation of losses when the investor does not neglect climate risk.
@@
1:Initialize LossLoss and $Loss_{shock}$